OpenStudy (anonymous):

lim x=0 (x+2)^3 - 8 / x

OpenStudy (anonymous):

is this an l'hopitals rule question?

sam (.sam.):

If you use L'hospital's rule you'll get \[\lim_{x \rightarrow 0}\frac{3(x+2)^2}{1} \\ \\ \lim_{x \rightarrow 0} 3(x+2)^2\] Solve

terenzreignz (terenzreignz):

But the numerator doesn't go to 0?

terenzreignz (terenzreignz):

Oh wait, yes it does... derp :3

OpenStudy (anonymous):

yes its a indetermination 0/0

terenzreignz (terenzreignz):

Well, @.Sam. pretty much clears it up :D

OpenStudy (anonymous):



OpenStudy (anonymous):

what was wrong with the answer @.Sam. gave?

OpenStudy (anonymous):

@MarilynRobles you factorized a bit wrong, if you're going to magically add 8 to one side, you need to add 8 to the other side. you have 18 instead

OpenStudy (anonymous):

thaknss

OpenStudy (anonymous):

why 8!! pleasee can you draw? its hard to me understand youu i dont speak english

OpenStudy (anonymous):

err sorry. x^3 + 18x -8 = 0 x^3 + 18x -8 +8 = +8 that was the step with a mistake

OpenStudy (anonymous):

and the x in the denominator? equation becomes?

OpenStudy (anonymous):

Use L'Hopital's rule to answer this problem, do not factorize when you plug in 0 and the equation is indeterminate, take the derivative of the numerator and denominator then plug in 0 again

sam (.sam.):

The mistake is the expansion of \((x+2)^3\) \[x^3+6 x^2+12 x+8\] \[\lim_{x \rightarrow 0}\frac{x^3+6 x^2+12 x+8-8}{x} \\ \\ \lim_{x \rightarrow 0}\frac{x^3+6 x^2+12 x}{x}\\ \\ \lim_{x \rightarrow 0}\frac{(x)(x^2+6 x+12) }{x} \\ \\ \lim_{x \rightarrow 0}x^2+6 x+12 \] Substitute x=0, 12

OpenStudy (anonymous):

thanks to all :)